OpenStudy (anonymous):

verify this equation cos^2x-sin^2x=1-2sin^2x

OpenStudy (anonymous):

using the base identity: \[sin^2(x)+cos^2(x)=1\] we can replace 1 with that and it should all simplify itsself out

OpenStudy (dumbcow):

use identity sin^2 + cos^2 = 1 --> cos^2 = 1 - sin^2 cos^2 - sin^2 = (1-sin^2) - sin^2 = 1-2sin^2

OpenStudy (anonymous):

yall are both quite helpful thanks